# SPVLoc: Semantic Panoramic Viewport Matching for 6D Camera Localization in Unseen Environments

### v1

@@ Abstract
In this paper, we present SPVLoc, a global indoor localization method that accurately determines the six-dimensional (6D) camera pose of a query image and requires minimal scene-specific prior knowledge and no scene-specific training. Our approach employs a novel matching procedure to localize the perspective camera’s viewport, given as an RGB image, within a set of panoramic semantic layout representations of the indoor environment. The panoramas are rendered from an untextured 3D reference model, which only comprises approximate structural information about room shapes, along with door and window annotations. We demonstrate that a straightforward convolutional network structure can successfully achieve image-to-panorama and ultimately image-to-model matching. Through a viewport classification score, we rank reference panoramas and select the best match for the query image. Then, a 6D relative pose is estimated between the chosen panorama and query image. Our experiments demonstrate that this approach not only efficiently bridges the domain gap but also generalizes well to previously unseen scenes that are not part of the training data. Moreover, it achieves superior localization accuracy compared to the state of the art methods and also estimates more degrees of freedom of the camera pose. We will make our source code publicly available at: https://github.com/fraunhoferhhi/spvloc.
@@ 1 Introduction
Recent methods incorporate synthetic data during training, but these localization models require retraining or fine-tuning for new scenes using detailed BIM (Building Information Modeling) models [1, 3]. The adoption of the Structured3D format [44] streamlines and standardizes 3D structure annotation by simplifying complex models into their most fundamental geometric representations. Methods that matching a query image to this simplified structural representation exhibit robustness in unexplored, dynamic environments [16, 28]. However, existing methods [28] mainly focus on 2D camera localization with respect to floor plans, missing two degrees of freedom for rotation and one degree of freedom for position. Also, some methods soley rely panoramic queries [16, 15]. These queries are captured under controlled conditions, involving a tripod-mounted camera with a precise horizontal alignment, and a fixed camera height [16]. However, in contexts such as AR applications or scenarios where archived photos must be accurately registered, e.g. to annotate visible building details into a digital building model, idealized capture conditions cannot always be guaranteed. In such cases, the demand for precise 6D localization becomes imperative.
In this paper, we present SPVLoc, a novel method for indoor 6D camera localization in unseen environments. It represents a 3D scene with a set of synthetically generated omnidirectional panorama images depicting fundamental semantic labels such as door, window, wall, floor, and ceiling. The core idea involves correlating the deep embedding of a perspective RGB camera image with the embeddings of panoramas to estimate the camera viewport – the precise region visible to the camera within the panorama. Leveraging a RetinaNet-like architecture [23], the 2D-bounding box around the viewport is predicted and classified. Retrieval of the closest panorama is performed by selecting the panorama with the globally highest classification score. The exact 6D pose is then found via relative pose estimation with an MLP starting from the panorama’s position. Our method offers several advantages. It relies solely on a minimalist 3D semantic building model [44], which could even be automatically extracted from simple 2D plans[8, 26, 32]. By bridging the domain gap between semantic representation and real images, it facilitates matching in unseen environments. Our approach learns wide baseline relative pose estimation and accurately predicts poses with few reference renderings.
@@ 2 Related Work
This section provides an overview of indoor camera localization and reviews related work on relative pose regression and perspective-to-panorama matching.
@@
Indoor camera localization is a longstanding challenge in the computer vision community and in recent years, with the advance of neural networks, several deep learning based approaches have been proposed. Typically, these methods rely on prior scene knowledge to built a scene representation and estimate a relation to a query image. Some methods use a 3D model generated via Structure-from-Motion [5] from registered camera images of the indoor environment and calculate the pose from 2D-3D correspondences, which might be either sparse [25, 41] or dense [36]. Alternatively, encoding the entire scene within a neural network and performing regression to determine the absolute pose is a frequently used strategy [29].
By leveraging synthetic renderings from a detailed 3D BIM model, Acharya et al. [1, 4] eliminated the need for real-world reference data. Their approach achieves real-time indoor localization with an accuracy of approximately 2 meters but is limited to the building of its origin, lacking generalizability to other structures. Liu et al. [24] pioneered floor plan-based image localization, predicting the facing wall within a cuboidal layout without estimating a full 3D pose. Howard et al. [16, 15] achieved 2D localization in unvisited scenes by aligning inferred room layouts with panoramic reference renderings [16] or encoded floor plans [15].
PF-net [18] addresses floor plan-based localization using a recurrent neural network optimized for a differentiable particle filter, designed for sequential updating but suboptimal for single images [28]. LASER [28] localizes the camera with respect to a floor plan, and considers both perspective and panoramic queries. It also estimates rotation around the yaw axis. The rendering code book scheme efficiently encodes floor plan information but is limited to 2D data, which constrains the network’s ability to estimate the remaining degrees of freedom.
To sum up, existing methods either depend on accurate scene specific data for a detailed 3D representation or do not estimate full 6D pose. In contrast, our method neither depends on real-world reference images, a detailed 3D model, or scene specific training but is still able to estimate the full 6D perspective camera pose by introducing a new cross-domain panoramic-to-perspective image registration approach between RGB images and semantic panoramas.
@@
Retrieval-based methods like NetVLAD [7] and DenseVLAD [38] start by identifying the most similar images in a database, providing a starting point for RPR. PixLoc [33] predicts multiscale deep features for a direct alignment from the reference and the query image with a CNN and use geometric optimization for calculating the pose. Turkoglu et al. [39] employ a GNN to re-localize cameras with respect to multiple frames. A graph links the query image to training counterparts, and a GNN refines representations for a consistent camera pose estimate. Our method shares similarities with [39] but goes further by addressing the domain gap between simplified synthetic and real data and incorporating multimodal matching between panoramic reference and perspective query images.
@@
Recently, perspective-to-panoramic matching advanced, due to sliding window approaches in feature space [30, 34]. It improved visual recognition compared to methods that simply sampled perspective reference images from panoramas [42]. Orhan et al. [31] explored leveraging semantic segmentation similarity to address appearance variations between reference and query images.
Our approach enhances perspective-to-panorama matching using virtual scene renderings as panoramic references. Inspired by multimodal zero-shot retrieval [17] and object detection, particularly Mercier et al.’s work [27], it utilizes the panorama as a global context for query matching.
@@ 3 Method
Figure 2: Network Architecture Overview: The information from the query branch is fed into the panorama branch by depth-wise correlation. Three task heads predict the corresponding viewport, one task head predicts the relative pose offset.
@@ 3.1 Semantic Panoramic Viewport Matching
Our objective is to precisely estimate the 6D camera pose denoted as $P$ for a given query RGB image $I_{q}$ relative to a semantic 3D model $S$. To achieve this, we reformulate the 6D indoor localization of $I_{q}$ within $S$ as a cross-domain image-to-panorama matching problem. The model is represented by a set of $N$ equirectangular reference layouts $L=\{L_{i}~{}|~{}i=1,\ldots N\}$ consisting of semantic labels in a single channel. Each rendering comprehensively captures the entire environment around a specific vantage point. The viewport $V$ is a distinct ’cutout’ section from a panorama, defining the exact region a camera positioned nearby can see. The perspective distortion of $V_{P_{i}}$ contains information about the relative pose offset $P_{i}$ between $L_{i}$ and $I_{q}$.
Scoring and ranking viewports helps in choosing the optimal panorama $L^{\star}$. The absolute pose $P$ is then formed by concatenating $P^{\star}$ with $P_{i}$. Identifying the viewport and the corresponding pose offset poses several challenges. First, with increasing distance between the two cameras the perspective distortion between image and panorama increases. Second, the matching involves a severe domain gap as a photographic RGB image is matched against a rendered semantic panorama.
@@
We create semantic panorama reference renderings utilizing a ray-tracing-based renderer [22]. The viewport mask $V_{P}$ is a binary image that signifies the visibility of points within the equirectangular image from the view of the perspective camera. The calculation involves projecting all 3D points from the panorama in the perspective view. The projection is performed based on the known underlying layout during the training phase. The viewport mask is determined by identifying pixels projecting into the perspective image, taking into account the 3D camera pose, and excluding those covered by occlusion.
@@
Our matching is inspired by the object detection strategy of DTOID [27]. Fig. 2 depicts an illustration of the network architecture. Initially, $I_{q}$ is encoded with an EfficientNet-S backbone. The features obtained after the fourth and the last downsampling layer are preserved. Subsequently, both outputs are rescaled to a tensor of spatial size of $7\times 7$ and concatenated to a tensor $F$.
@@ 3.2 Feature-Correlation-based Pose Regression
An additional Pose head processes $R^{\star}$ to estimate the relative pose offset of the camera concerning a panorama rendering (Fig. 3). We first utilize three convolution blocks, whereby the first two are follwed by a max-pooling layer, to compress the features. The flattened features form an 1D embedding $r^{\star}\in\mathbb{R}^{C}$. Here, $C=640$ for a panorama of dimensions $256\times 128$. Optionally, to handle different camera field of views during training and testing, $r^{\star}$ is normalized and multiplied by the horizontal field of view angle in radians. The result is fed into a four-layer MLP, producing the 3D offset to the query image location $\hat{t}{i}\in\mathbb{R}^{3}$ and a rotation given as the logarithm of the unit quaternion $\hat{r}{i}=\log(\hat{q})_{ij}\in\mathbb{R}^{3}$ [39].
@@ 3.4 Inference
For pose estimation, we select reference positions with the top-$n$ highest bounding box classification scores. The absolute pose is determined by the result of the Pose head, where the selected reference position is added to the translation. To eliminate the necessity for rendering during inference, we propose pre-generating reference panoramas for a floor plan. Additionally, the encoding of these panoramas can be precomputed. This ensures that only the encoded feature $R$, need to be stored, significantly streamlining the inference process.
@@ 4 Experiments
consists of 3500 near-photorealistic professionally modeled indoor environments, each annotated with ground truth 3D structure information. It encompasses 21835 panoramic images, each provided in both furnished and unfurnished configurations. The authors define a unified representation, describing the scene with geometric plane, line, and junction point primitives, their incidences and intersections, as well as semantic annotations. The semantic classes include floor, ceiling, wall, door, and window.
@@
comprises 67,448 panorama images taken in 1,575 unfurnished residential homes, all globally aligned and registered to a floor plan. The 3D annotations of panoramas within a scene are adequate for transforming them into the S3D representation format. Before training, we transform all data to that format. In contrast to S3D, the annotation includes openings between rooms, e.g. if a corridor is linked to a room without a door. These openings increase the structural variation of the semantic renderings. Fig. 4 shows an example reference model, generated from Zillow annotations.
@@
For both datasets, we employ the official train-test scene split. To enhance control over the field of view and angle variation, we sample perspective test images from the RGB panoramas. They are registered in the 3D model, so the semantics are known for each perspective image. We compile test-sets suitable for 6D pose estimation and structured evaluation  [^1] Our code, including scripts to reproduce the test data will be published online.. The test sets consist of square perspective images with a viewing angle of $60$, $90$, and $120^{\circ}$, respectively, along with variations in roll and pitch, and random yaw angles for ZinD and S3D. To ensure meaningful views, we apply two criteria on the sampled yaw. First, the images contain a minimum of three semantic classes, like wall-floor-ceiling or wall-window-floor. Second, the camera faces inside the room rather than towards the next wall. The criteria are applied to the $60^{\circ}$ images (4420 for ZiND and 1133 for S3D). Subsequent variations in opening angles and roll-pitch configurations are generated from this starting point.
@@ 4.1 Details
For evaluation, we report the fraction of frames localized correctly within a threshold $\tau$. We also list the median rotation and translation error, as well as the fraction of frames localized within an additional angle threshold $\upsilon$ and $\tau$. In the retrieval process, the top-3 best matches are selected, and we report recall metrics for both top-1 and the best among the top-3.
@@
For the state of the art comparison, we train a model with a variable field of view, while other experiments use a model trained with a fixed opening angle of $90^{\circ}$. The model with variable viewing angles randomly samples views between $45$ and $135^{\circ}$ during training.
We consistently render $s$ randomly positioned panoramas around the query within a radius of ±$\pm r_{1}$ in $xy$ direction and ±$\pm r_{2}$ in $z$ (upward) direction. For a single query, the viewport and relative pose offset is therefore calculated against multiple references. In our experimental setup, we set $s$ to 4, $r_{1}$ to 1.4 meters, and $r_{2}$ to 30 centimeters. Additionally, one random negative example is generated in a different room with the target to enhance the network’s ability to distinguish subtle room differences. Images are sampled with random yaw and ±$\pm 10^{\circ}$ random pitch and roll angle. The batch size is set to 40, comprising 40 query images and 200 panoramas, and the training is performed on a single NVIDIA A100 GPU. Query images with fewer than three semantic classes are disregarded during the loss calculation process. Training runs for approximately 42,000 steps with an initial learning rate of $2.5\times 10^{-4}$, halved two times during training.

### v2

@@ Abstract
In this paper, we present SPVLoc, a global indoor localization method that accurately determines the six-dimensional (6D) camera pose of a query image and requires minimal scene-specific prior knowledge and no scene-specific training. Our approach employs a novel matching procedure to localize the perspective camera’s viewport, given as an RGB image, within a set of panoramic semantic layout representations of the indoor environment. The panoramas are rendered from an untextured 3D reference model, which only comprises approximate structural information about room shapes, along with door and window annotations. We demonstrate that a straightforward convolutional network structure can successfully achieve image-to-panorama and ultimately image-to-model matching. Through a viewport classification score, we rank reference panoramas and select the best match for the query image. Then, a 6D relative pose is estimated between the chosen panorama and query image. Our experiments demonstrate that this approach not only efficiently bridges the domain gap but also generalizes well to previously unseen scenes that are not part of the training data. Moreover, it achieves superior localization accuracy compared to the state of the art methods and also estimates more degrees of freedom of the camera pose. Our source code is publicly available at: https://fraunhoferhhi.github.io/spvloc.
@@ 1 Introduction
Recent methods incorporate synthetic data during training, but these localization models require retraining or fine-tuning for new scenes using detailed BIM (Building Information Modeling) models [1, 3]. The adoption of the Structured3D format [46] streamlines and standardizes 3D structure annotation by simplifying complex models into their most fundamental geometric representations. Methods that match a query image to this simplified structural representation exhibit robustness in unexplored, dynamic environments [17, 30]. However, existing methods [30, 9] mainly focus on 2D perspective camera localization with respect to floor plans, missing two degrees of freedom for rotation and one degree of freedom for position. Also, some methods solely rely on panoramic queries [17, 16, 23]. These queries are typically captured under controlled conditions, involving a tripod-mounted camera with a precise horizontal alignment, and a fixed camera height [17]. However, in AR applications or scenarios where archived photos must be accurately registered, e.g. to annotate visible building details into a digital building model, idealized capture conditions cannot always be guaranteed. In such cases, the demand for precise 6D localization becomes imperative.
In this paper, we present SPVLoc, a novel method for indoor 6D camera localization in unseen environments. It represents a 3D scene with a set of synthetically generated omnidirectional panorama images depicting fundamental semantic labels such as door, window, wall, floor, and ceiling. The core idea involves correlating the deep embedding of a perspective RGB camera image with the embeddings of panoramas to estimate the camera viewport – the precise region visible to the camera within the panorama. Leveraging a RetinaNet-like architecture [25], the 2D-bounding box around the viewport is predicted and classified. Retrieval of the closest panorama is performed by selecting the panorama with the globally highest classification score. The exact 6D pose is then found via relative pose estimation with an MLP starting from the panorama’s position. Our method offers several advantages. It relies solely on a minimalist 3D semantic building model [46], which could even be automatically extracted from simple 2D plans[8, 28, 34]. By bridging the domain gap between semantic representation and real images, it facilitates matching in unseen environments. Our approach (Fig. 1) learns wide baseline relative pose estimation and accurately predicts poses with few reference renderings.
@@ 2 Related Work
Indoor camera localization is a longstanding challenge in the computer vision community and in recent years, with the advance of neural networks, several deep learning based approaches have been proposed. Typically, these methods rely on prior scene acquisition to built a detailed scene representation and estimate a relation to a query image. Some methods use a 3D model generated via Structure-from-Motion [5] from registered camera images of the indoor environment and calculate the pose from 2D-3D correspondences, which might be either sparse [27, 43] or dense [38]. Alternatively, encoding the entire scene within a neural network and performing regression to determine the absolute pose is a frequently used strategy [31].
By leveraging synthetic renderings from a detailed 3D BIM model, Acharya et al. eliminated the need for real-world reference data. Their approaches achieve real-time indoor localization [1] and an accuracy of approximately 1 meter [4] but are limited to the building of their origin, lacking generalizability to other structures. Liu et al. [26] pioneered floor plan-based image localization, predicting the facing wall within a cuboidal layout without estimating a full 3D pose. Howard et al. achieved 2D panorama localization in unseen scenes by aligning inferred room layouts with panoramic reference renderings [17] or encoded floor plans [16].
Several methods estimate the 2D position and 1D rotation of a perspective camera relative to a floor plan. PF-net [19] uses a recurrent neural network optimized for a differentiable particle filter, designed for sequential updating but suboptimal for single images [30]. F3Loc [9] improves sequential 2D localization against a floorplan by integrating multiview geometry cues. LASER [30] localizes the camera with respect to a semantic floor plan, and considers both perspective and panoramic queries. The rendering code book scheme efficiently encodes floor plan information but is limited to 2D data, which constrains the network’s ability to estimate the remaining degrees of freedom.
To sum up, existing methods either depend on accurate scene specific data for a detailed 3D representation or do not estimate the 6D camera pose. In contrast, our method neither depends on real-world reference images, a detailed 3D model, or scene specific training but is still able to estimate the 6D perspective camera pose by introducing a new cross-domain panoramic-to-perspective image registration approach between RGB images and semantic panoramas.
@@
Retrieval-based methods like NetVLAD [7] and DenseVLAD [40] start by identifying the most similar images in a database, providing a starting point for RPR. PixLoc [35] predicts multiscale deep features for direct alignment of reference and query image with a CNN and uses geometric optimization for calculating the pose. Turkoglu et al. [41] employ a GNN to re-localize cameras with respect to multiple frames. A graph links the query image to training counterparts, and a GNN refines representations for a consistent camera pose estimate. Our method shares similarities with [41] but goes further by addressing the domain gap between simplified synthetic and real data, and by incorporating multimodal matching between panoramic reference and perspective query images.
@@
Recently, perspective-to-panoramic matching advanced, due to sliding window approaches in feature space [32, 36]. It improved visual recognition compared to methods that simply sampled perspective reference images from panoramas [44]. Orhan et al. [33] explored leveraging semantic segmentation similarity to address appearance variations between reference and query images. Our approach enhances perspective-to-panorama matching using virtual scene renderings as panoramic references. Inspired by multimodal zero-shot retrieval [18] and object detection, particularly Mercier et al.’s work [29], it utilizes the panorama as a global context for query matching.
@@ 3 Method
Figure 2: Network architecture overview. The information from the query branch is fed into the panorama branch by depth-wise correlation. Three task heads predict the corresponding viewport, one task head predicts the relative pose offset.
@@ 3.1 Semantic Panoramic Viewport Matching
Our objective is to precisely estimate the 6D camera pose denoted as $P$ for a given query RGB image $I_{q}$ relative to a semantic 3D model $S$. To achieve this, we reformulate the 6D indoor localization of $I_{q}$ within $S$ as a cross-domain image-to-panorama matching problem. The model is represented by a set of $N$ equirectangular reference layouts $L=\{L_{i}~{}|~{}i=1,\ldots N\}$ consisting of semantic labels in a single channel. Each rendering comprehensively captures the entire environment around a specific vantage point. The viewport $V$ is a distinct cutout section from a panorama, defining the exact region a camera positioned nearby can see. The perspective distortion of $V_{P_{i}}$ contains information about the relative pose offset $P_{i}$ between $L_{i}$ and $I_{q}$. Scoring and ranking viewports helps in choosing the optimal panorama $L^{\star}$. The absolute pose $P$ is then formed by concatenating $P^{\star}$ with $P_{i}$. Identifying the viewport and the corresponding pose offset poses several challenges. First, with increasing distance between the two cameras the perspective distortion between image and panorama increases. Second, the matching involves a severe domain gap as a photographic RGB image is matched against a rendered semantic panorama.
@@
Semantic panorama reference renderings are created utilizing a ray-tracing-based renderer [24]. The viewport mask $V_{P}$ is a binary image that signifies the visibility of points within the equirectangular image from the view of the perspective camera. The calculation involves projecting all 3D points from the panorama in the perspective view. The projection is performed based on the known underlying layout during the training phase. The mask $V_{P}$ is determined by identifying pixels projecting into the perspective image, taking into account the 3D camera pose, and excluding those covered by occlusion.
@@
Our matching is inspired by the object detection strategy of DTOID [29]. Figure 2 illustrates the network architecture. Initially, $I_{q}$ is encoded with an EfficientNet-S backbone. The features obtained after the fourth and the last downsampling layer are preserved. Subsequently, both outputs are rescaled to a tensor of spatial size of $7\times 7$ and concatenated to a tensor $F$.
@@ 3.2 Feature-Correlation-based Pose Regression
An additional Pose head processes $R^{\star}$ to estimate the relative pose offset of the camera concerning a panorama rendering (Fig. 3). We first utilize three convolution blocks, whereby the first two are followed by a max-pooling layer, to compress the features. The flattened features form an 1D embedding $r^{\star}\in\mathbb{R}^{C}$. Here, $C=640$ for a panorama of dimensions $256\times 128$. Optionally, to handle different camera field of views during training and testing, $r^{\star}$ is normalized and multiplied by the horizontal field of view angle in radians. The result is fed into a four-layer MLP, producing the 3D offset to the query image location $\hat{t}{i}\in\mathbb{R}^{3}$ and a rotation given as the logarithm of the unit quaternion $\hat{r}{i}=\log(\hat{q})_{ij}\in\mathbb{R}^{3}$ [41].
@@ 3.4 Inference
For pose estimation, we select reference positions with the top-$n$ highest bounding box classification scores. The absolute pose is determined by the result of the Pose head, where the selected reference position is added to the translation. To eliminate the necessity for rendering during inference, we propose pre-generating reference panoramas for a floor plan. Additionally, the encoding of these panoramas can be precomputed. This ensures that only the encoded features $R$ need to be stored, significantly streamlining the inference process.
@@ 4 Experiments
consists of 3500 near-photorealistic professionally modeled indoor environments, each annotated with ground truth 3D structure information. It encompasses 21835 panoramic images, each provided in both furnished and unfurnished configurations. The authors define a unified representation that describes scenes using geometric primitives such as planes, lines, junction points, and semantic annotations including floor, ceiling, wall, door, and window classes.
@@
includes 67,448 panorama images from 1,575 unfurnished residential homes, globally aligned and registered to floor plans. The 3D annotations of panoramas within a scene are adequate for transforming them into the S3D representation format. Before training, we transform all data to that format. Figure 4 shows an example reference model generated from ZInD annotations. Unlike S3D, ZInD includes room openings such as corridors connected to rooms without doors, enhancing structural variation in semantic renderings.
@@
For both datasets, we employ the official train-test scene split. To enhance control over the field of view and angle variation, we sample perspective test images from the RGB panoramas. They are registered in the 3D model, so the semantics are known for each perspective image. We compile test-sets suitable for 6D pose estimation and structured evaluation. The test sets consist of square perspective images with a viewing angle of $60$, $90$, and $120^{\circ}$, respectively, along with variations in roll and pitch, and random yaw angles for ZinD and S3D. To ensure meaningful views, we apply two criteria on the sampled yaw. First, the images contain a minimum of three semantic classes, like wall-floor-ceiling or wall-window-floor. Second, the camera faces inside the room rather than towards the next wall. The criteria are applied to the $60^{\circ}$ images (4420 for ZiND and 1133 for S3D). Subsequent variations in opening angles and roll-pitch configurations are generated from this starting point.
@@ 4.1 Details
For evaluation, we report the fraction of frames localized correctly within a threshold $\tau$. We also list the median rotation and translation error, as well as the fraction of frames localized within $\tau$ and an additional angle threshold $\upsilon$. In the retrieval process, the top-3 best matches are selected, and we report recall metrics for both top-1 and the best among the top-3.
@@
For the state of the art comparison, we train a model with a variable field of view, while other experiments use a model trained with a fixed opening angle of $90^{\circ}$. The model with variable viewing angles randomly samples views between $45$ and $135^{\circ}$ during training. We consistently render $s$ randomly positioned panoramas around the query within a radius of ±$\pm r_{1}$ in $xy$ direction and ±$\pm r_{2}$ in $z$ (upward) direction. For a single query, the viewport and relative pose offset is therefore calculated against multiple references. In our experimental setup, we set $s$ to 4, $r_{1}$ to 1.4 meters, and $r_{2}$ to 30 centimeters. Additionally, one random negative example is generated outside $r_{1}$, likely in a different room, with the target to enhance the network’s ability to distinguish subtle room differences. Images are sampled with random yaw and ±$\pm 10^{\circ}$ random pitch and roll angle. The batch size is set to 40, comprising 40 query images and 200 panoramas, and the training is performed on a single NVIDIA A100 GPU. Query images with fewer than three semantic classes are disregarded during the loss calculation process. Training runs for approximately 42,000 steps with an initial learning rate of $2.5\times 10^{-4}$, halved two times during training.